Find two numbers such that each of the
numbers is the square of the other number.
The two numbers should NOT be equal!

Answer by Alan3354(69443)

y = x^2
x = y^2
------------
x = (x^2)^2 = x^4
x^4 - x = 0
Only 1 and zero fit, and in both cases x = y.